OpenStudy (anonymous):

which is bigger 2/3 or 3/5....

OpenStudy (diyadiya):

2/3

OpenStudy (diyadiya):

2/3=0.666.. 3/5=0.6

OpenStudy (anonymous):

haha...thnks ma lttle could was lyk i want proof of tht! lol now i can show him i wuz rite!!!!

OpenStudy (dumbcow):

Find common denominator so you can compare........2/3 = 10/15.....3/5 = 9/15

OpenStudy (anonymous):

hey